OpenStudy (anonymous):

Convert x = (3pi)/4 to rectangular form.

ganeshie8 (ganeshie8):

you sure it is "x" ? and its not \(\theta\) ?

OpenStudy (anonymous):

oops yes its theta

ganeshie8 (ganeshie8):

\(\theta = \dfrac{3\pi}{4}\) like this right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

use below : \(\tan \theta = \dfrac{y}{x}\)

ganeshie8 (ganeshie8):

\(\tan (\frac{3\pi}{4}) = \dfrac{y}{x}\)

ganeshie8 (ganeshie8):

\(-1 = \dfrac{y}{x}\) \(y = -x\)

OpenStudy (anonymous):

oh ok how did you get -1?

ganeshie8 (ganeshie8):

\(\tan (\frac{3\pi}{4}) = \tan ( \pi - \frac{\pi}{4}) = -\tan(\frac{\pi}{4}) = -1\)